# SVEMnet: An R package for Self-Validated Elastic-Net Ensembles and Multi-Response Optimization in Small-Sample Mixture–Process Experiments

### v2

@@ Abstract
SVEMnet is an R package for fitting Self-Validated Ensemble Models (SVEM) with elastic-net base learners and performing multi-response optimization in small-sample mixture–process design-of-experiments (DOE) studies with numeric, categorical, and mixture factors. SVEMnet wraps elastic-net and relaxed elastic-net models for Gaussian and binomial responses from glmnet in a fractional random-weight (FRW) resampling scheme with anti-correlated train/validation weights; penalties are selected by validation-weighted AIC- and BIC-type criteria, and predictions are averaged across replicates to stabilize fits near the interpolation boundary. In addition to the core SVEM engine, the package provides deterministic high-order formula expansion, a permutation-based whole-model test heuristic, and a mixture-constrained random-search optimizer that combines Derringer–Suich desirability functions, bootstrap-based uncertainty summaries, and optional mean-level specification-limit probabilities to generate scored candidate tables and diverse exploitation and exploration medoids for sequential fit–score–run–refit workflows. A simulated lipid nanoparticle (LNP) formulation study illustrates these tools in a small-sample mixture–process DOE setting, and simulation experiments based on sparse quadratic response surfaces benchmark SVEMnet against repeated cross-validated elastic-net baselines.
@@ 1 Introduction
This paper introduces SVEMnet, an R package that implements SVEM with elastic-net and optional relaxed elastic-net base learners [8, 9, 10, 11]. In a relaxed fit, glmnet computes a penalized solution path indexed by the penalty parameter $\lambda$, identifies the active set at each $\lambda$, and then refits an unpenalized linear model on that active set, interpolating between penalized and unpenalized coefficients via a relaxation parameter $\gamma\in[0,1]$. In the sparse Gaussian linear model studied by Meinshausen [10], the relaxed lasso reduces the shrinkage of non-zero coefficients and can achieve lower prediction error than the standard lasso. Simulations in Section 4 indicate that combining relaxed base learners with the SVEM FRW ensemble can further reduce out-of-sample prediction error for Gaussian responses, plausibly because the ensemble averages over the additional variance introduced by relaxation.
@@
In place of the validation-weighted sum of squared errors criterion (denoted wSSE) used in earlier SVEM work by Lemkus et al. [2], SVEMnet standardizes on heuristic, validation-weighted AIC/BIC-like criteria—wAIC for Gaussian responses and wBIC for binomial responses [12, 13, 14]. Flexible fits such as the lasso are known to behave poorly near the interpolation boundary, where the number of runs $n$ approaches the full expansion dimension $p_{\mathrm{full}}$ (the total number of coefficients in the “full” linear model expansion, including the intercept), with “peaking” in out-of-sample error documented in small-sample classification and regression problems [15] and in lasso model selection [16]. In the simulations reported in Section 4, wAIC and wBIC avoid this peaking, whereas the loss-only wSSE selector exhibits a pronounced spike at $n\approx p_{\mathrm{full}}$.
@@
Self‑Validated Ensemble Models have already seen practical use in industrial DOE and chemometric settings, including mixture–process workflows for lipid nanoparticle (LNP) formulations and other pharmaceutical and analytical applications [17, 18, 6, 19, 20, 21, 22, 23, 24, 25]. These applications motivate an implementation tailored to chemometric DOE. In particular, SVEMnet is designed to automate many of the steps in the SVEM-based mixture–process workflow for LNP formulation optimization described by Karl et al. [17]. SVEMnet supplies an open-source SVEM implementation over glmnet with validation-weighted AIC/BIC selectors, deterministic high-order full model expansions, a permutation-based whole-model test heuristic, and a mixture-constrained random-search scoring tool for multi-response desirability and candidate generation.
@@
SVEMnet is intended for round-by-round iteration in small-$N$ DOE: fit SVEM models on the current data; use the random-search optimizer to generate and score a feasible candidate set under mixture and process constraints; select a small number of high-score exploitation medoids and high-uncertainty exploration medoids; then run those settings, append the new data, and refit the models. The lipid nanoparticle (LNP) formulation example below illustrates this workflow [17, 5]. The sequential workflow is an iterate–append loop that alternates exploitation and exploration based on predictive uncertainty, rather than a full Gaussian-process-based Bayesian optimization scheme, which remains the standard choice when richer covariance structures and acquisition strategies are desired [26, 27].
@@
SVEMnet can be used with any single-error-stratum experimental design in which individual runs can be treated as independent. However, the predictive performance of SVEMnet will naturally depend on design quality. For physical experiments we recommend optimal designs [28, 29], fast flexible space-filling designs [30], or U-Bridge designs [31] rather than the simple Latin hypercube samples [32] that are employed for computational expediency in the simulations in Section 4. Because the FRW bootstrap treats runs as the resampling units, split-plot and other multi-stratum designs with hierarchical variance components are not currently accommodated; blocking factors may only be treated as fixed effects.
@@
The remainder of the paper is organized as follows. Section 2 details the SVEMnet modeling engine: SVEM background, the glmnet wrapper (including relaxed refits), deterministic expansion tools, the validation-weighted wAIC/wBIC selector, and the multi-response random-search optimizer. Section 3 illustrates the workflow on a lipid nanoparticle (LNP) formulation example. Section 4 reports simulations, and the paper concludes with a summary and outlook. Technical specifics of the validation-weighted criterion and mixed-model multiple-comparison tables for simulation results appear in A and B.
@@ 2.2 SVEM background
$w_{i}^{\mathrm{train}}=-\log U_{i},\qquad w_{i}^{\mathrm{valid}}=-\log(1-U_{i}),$
@@
Given $(w^{\mathrm{train}},w^{\mathrm{valid}})$ for replicate $b$, SVEMnet fits an elastic-net or relaxed elastic-net base learner to the FRW-weighted training sample under $w^{\mathrm{train}}$, evaluates a collection of path points using a validation-weighted information criterion computed with $w^{\mathrm{valid}}$ (Section 2.4), and stores predictions from the selected path point. Overall SVEM predictions are obtained by averaging these member predictions across $b=1,\dots,B$. Lemkus et al. [2] tuned penalties by validation-weighted SSE (wSSE); SVEMnet adopts the same FRW-and-ensemble structure but standardizes on validation-weighted AIC/BIC analogs—wAIC default for Gaussian responses; wBIC default for binomial responses.
Although we focus on Gaussian responses in the exposition and case study, the software also implements a binomial (logistic) variant. In that case, the elastic-net base learner is fit under an FRW-weighted logistic log-likelihood, penalties are selected on the deviance scale, and ensemble predictions are formed by averaging member-predicted probabilities; A gives the corresponding criteria.
@@ 2.3.1 Deterministic expansion utilities
A practical convenience in SVEMnet is that the user lists main-effect predictors only once and lets the expansion utilities (bigexp_terms() and bigexp_formula()) deterministically build interaction and polynomial columns. The same specification can be reused across responses, guaranteeing identical design matrices, stable column order, and reproducible comparisons. With factorial_order = 2 and polynomial_order = 2, for example, the generated right-hand side includes all two-way interactions and quadratic terms (and, optionally, partial-cubic terms of the form $\mathrm{I}(X^{2})\!:\!Z$) without hand-writing products or powers; categorical factors are handled using the contrasts recorded for the training data. When blocking or covariate factors (e.g., Operator, day, plate, ambient temperature) are supplied via blocking in bigexp_terms(), these factors make an additive adjustment to the response surface and do not interact with the other study factors.
@@ 2.5.1 Sampling and specification limits
Candidate settings are generated by svem_random_table_multi(), which samples feasible combinations under any supplied mixture groups (sum-to-constant constraints with bounds) together with bounds on process variables taken from the factor ranges in the training data. When specification limits are provided for a response, SVEMnet uses the SVEM bootstrap to estimate, at each candidate, the probability that the process mean lies inside those limits and adds per-response and joint mean-level specification probabilities (e.g., prob_in_spec and p_joint_mean) as additional columns in the scored table. These mean-level probabilities are used only as ranking proxies: formal specification limits apply to individual observations, not to the process mean.
@@ 2.5.2 Scoring and response weighting
$S(x)=\exp\!\left(\sum_{r}w_{r}\log\big(d_{r,\varepsilon}(x)\big)\right),\qquad d_{r,\varepsilon}(x)=(1-\varepsilon)\,d_{r}(x)+\varepsilon,$
@@
User-specified goal weights $\tilde{w}_{r}$ are first normalized to $\bar{w}_{r}$ with $\sum_{r}\bar{w}_{r}=1$. When reweight_by_wmt = FALSE, we set $w_{r}=\bar{w}_{r}$ and report a single column score based on $S(x)$. When reweight_by_wmt = TRUE for Gaussian responses, we further reweight the normalized user weights by a monotone function of the permutation whole-model $p$-values from [5] (for example $\log_{10}(p_{r})$), renormalize to obtain WMT-adjusted weights $w_{r}^{\mathrm{WMT}}$, and compute a second geometric-mean score wmt_score by using $w_{r}=w_{r}^{\mathrm{WMT}}$ in $S(x)$. Methodologically, WMT compares the pattern of SVEM predictions on a space-filling set of design points with a reference distribution obtained by refitting the same SVEM specification to permuted responses, yielding an approximate whole-model $p$-value for the global null of a flat response surface [5]. These WMT-adjusted weights are diagnostic and do not alter the underlying SVEM fits. In this paper we treat the resulting WMT-based multipliers as a secondary diagnostic overlay: the primary optimization uses the user-specified weights $\tilde{w}_{r}$, and wmt_score is intended for sensitivity checks rather than as the sole ranking criterion.
@@
For candidate scoring, blocking variables are held fixed at a single reference setting while the controllable factors are varied over the feasible region. Categorical blocking factors are evaluated at the most common observed level; continuous blocking covariates are evaluated at the midpoint of their modeled range. Under the additive-only blocking assumption this induces only an additive offset in the predicted responses and therefore does not change the ranking of candidate recipes.

### v3

@@ Abstract
SVEMnet is an R package for fitting Self-Validated Ensemble Models (SVEM) with elastic-net base learners and performing multi-response optimization in small-sample mixture–process design-of-experiments (DOE) studies with numeric, categorical, and mixture factors. SVEMnet wraps elastic-net and relaxed elastic-net models for Gaussian and binomial responses from glmnet in a fractional random-weight (FRW) resampling scheme with anti-correlated train/validation weights; penalties are selected by validation-weighted AIC- and BIC-type criteria, and predictions are averaged across replicates to stabilize fits near the interpolation boundary. In addition to the core SVEM engine, the package provides deterministic high-order formula expansion, a permutation-based whole-model test heuristic, and a mixture-constrained random-search optimizer that combines Derringer–Suich desirability functions, bootstrap-based uncertainty summaries, and optional mean-level specification-limit probabilities to generate scored candidate tables and diverse exploitation and exploration medoids for sequential fit–score–run–refit workflows. A simulated lipid nanoparticle (LNP) formulation study illustrates these tools, and simulation experiments based on sparse quadratic response surfaces benchmark SVEMnet against repeated cross-validated elastic-net baselines.
@@
journal: Chemometrics and Intelligent Laboratory Systems
[…]
[inst1] organization=Karl Statistical Services, city=Aurora, state=CO, postcode=80016, country=USA
@@ 1 Introduction
SVEM has already been implemented in JMP Pro (with forward-selection and Lasso base learners) and a proprietary JMP add-in (using neural network base learners, [6]). However, many analysts work in R, and reproducible small-sample workflows benefit from an open-source implementation that can be scripted, versioned, and embedded in simulation and optimization pipelines. To this end, this paper introduces SVEMnet, an R package that implements SVEM with elastic-net and optional relaxed elastic-net base learners [8, 9, 10, 11]. In a relaxed fit, glmnet computes a penalized solution path indexed by the penalty parameter $\lambda$, identifies the active set at each $\lambda$, and then refits an unpenalized linear model on that active set, interpolating between penalized and unpenalized coefficients via a relaxation parameter $\gamma\in[0,1]$. In the sparse Gaussian linear model studied by Meinshausen [10], the relaxed lasso reduces the shrinkage of non-zero coefficients and can achieve lower prediction error than the standard lasso. Simulations in Section 4 indicate that combining relaxed base learners with the SVEM FRW ensemble can further reduce out-of-sample prediction error for Gaussian responses, plausibly because the ensemble averages over the additional variance introduced by relaxation. In place of the validation-weighted sum of squared errors criterion (denoted wSSE) used in earlier SVEM work by Lemkus et al. [2], SVEMnet standardizes on heuristic, validation-weighted information-criterion analogs—a validation-weighted Akaike information criterion (AIC) analog (wAIC) and a validation-weighted Bayesian information criterion (BIC; Schwarz criterion) analog (wBIC), with wAIC the default for Gaussian responses and wBIC the default for binomial responses [12, 13, 14]. Flexible fits such as the lasso are known to behave poorly near the interpolation boundary, where the number of runs $n$ approaches the full expansion dimension $p_{\mathrm{full}}$ (the total number of coefficients in the “full” linear model expansion, including the intercept), with “peaking” in out-of-sample error documented in small-sample classification and regression problems [15] and in lasso model selection [16]. Related analyses of the $p\gtrsim n$ interpolation setting in linear prediction underscore that performance can shift sharply near this boundary [17, 18]. In the simulations reported in Section 4, wAIC and wBIC avoid this peaking, whereas the loss-only wSSE selector exhibits a pronounced spike at $n\approx p_{\mathrm{full}}$.
@@
Self-Validated Ensemble Models have already seen practical use in industrial DOE and chemometric settings, including mixture–process workflows for lipid nanoparticle (LNP) formulations and other pharmaceutical and analytical applications [19, 20, 6, 21, 22, 23, 24, 25, 26, 27]. These applications motivate an implementation tailored to chemometric DOE. In particular, SVEMnet is designed to automate many of the steps in the SVEM-based mixture–process workflow for LNP formulation optimization described by Karl et al. [19]. SVEMnet supplies an open-source SVEM implementation over glmnet with validation-weighted AIC/BIC selectors, deterministic high-order full model expansions, a permutation-based whole-model test heuristic, and a mixture-constrained random-search scoring tool for multi-response desirability and candidate generation.
@@
SVEMnet is intended for round-by-round iteration in small-$N$ DOE: fit SVEM models on the current data; use the random-search optimizer to generate and score a feasible candidate set under mixture and process constraints; select a small number of high-score exploitation medoids and high-uncertainty exploration medoids; then run those settings, append the new data, and refit the models. The LNP formulation example below illustrates this workflow [19, 5]. The sequential workflow is an iterate–append loop that alternates exploitation and exploration based on predictive uncertainty, rather than a full Gaussian-process-based Bayesian optimization scheme, which remains the standard choice when richer covariance structures and acquisition strategies are desired [28, 29].
@@
SVEMnet can be used with any single-error-stratum experimental design in which individual runs can be treated as independent. However, the predictive performance of SVEMnet will naturally depend on design quality. For physical experiments we recommend optimal designs [30, 31], fast flexible space-filling designs [32], or U-Bridge designs [33] rather than the simple Latin hypercube samples [34] that are employed for computational expediency in the baseline simulations in Section 4. Because the FRW bootstrap treats experimental runs as the resampling units, SVEM is inherently a single-stratum procedure: split-plot and other multi-stratum designs with hierarchical variance components are not supported and would require substantial structural changes (e.g., hierarchical base learners and a different resampling unit). In SVEMnet, blocking or batch indicators may be included only as fixed additive effects.
@@
The remainder of the paper is organized as follows. Section 2 details the SVEMnet modeling engine: SVEM background, the glmnet wrapper (including relaxed refits), deterministic expansion tools, the validation-weighted wAIC/wBIC selector, and the multi-response random-search optimizer. Section 3 illustrates the workflow on an LNP formulation example. Section 4 reports simulations, and the paper concludes with a summary and outlook. Technical specifics of the validation-weighted criterion appear in A. Supplementary robustness checks and diagnostics appear in B. Mixed-model multiple-comparison tables for simulation results appear in C.
@@ 2.1 Software overview
    svem_random_table_multi(): generate feasible candidate settings under mixture and process constraints (candidate generation only);
  • 5.
@@
2.1.1 Workflow overview and typical use
The within-replicate tuning procedure used to select $(\alpha,\gamma,\lambda)$ is summarized in Figure 1 (Section 2.4). SVEMnet embeds this core engine in a sequential loop: (i) specify the factor space and construct a deterministic expansion (optional); (ii) fit SVEM models to one or more responses; (iii) optionally run the permutation-based whole-model diagnostic (WMT; Section 2.5) to assess global signal; (iv) generate a feasible candidate set under mixture and process constraints; (v) score candidates via multi-response desirability and summarize high-score (exploitation) and high-uncertainty (exploration) settings using medoids; and (vi) run selected settings, append the new rows, and refit. The optional wmt_score column is treated as a diagnostic overlay rather than the primary optimization objective.
@@ 2.2 SVEM background
$w_{i}^{\mathrm{train}}=-\log U_{i},\qquad w_{i}^{\mathrm{valid}}=-\log(1-U_{i}),$ (1)
@@
Given $(w^{\mathrm{train}},w^{\mathrm{valid}})$ for replicate $b$, SVEMnet fits an elastic-net or relaxed elastic-net base learner to the FRW-weighted training sample under $w^{\mathrm{train}}$, evaluates a collection of path points using a validation-weighted information criterion computed with $w^{\mathrm{valid}}$ (Section 2.4), and stores predictions from the selected path point. Overall SVEM predictions are obtained by averaging these bootstrap predictions across $b=1,\dots,B$. Lemkus et al. [2] tuned penalties by validation-weighted SSE (wSSE); SVEMnet adopts the same FRW-and-ensemble structure but standardizes on validation-weighted AIC/BIC analogs—wAIC default for Gaussian responses; wBIC default for binomial responses.
Although we focus on Gaussian responses in the exposition and case study, the software also implements a binomial (logistic) variant. In that case, the elastic-net base learner is fit under an FRW-weighted logistic log-likelihood, penalties are selected on the deviance scale, and ensemble predictions are formed by averaging bootstrap-member-predicted probabilities; A gives the corresponding criteria.
@@ 2.3 Implementation note: a SVEM wrapper over glmnet
Relaxation alters the usual bias–variance trade-off by reducing shrinkage after variable selection: individual relaxed fits can have lower bias but higher variance, especially near the interpolation boundary. In a SVEM ensemble, averaging across FRW replicates can partially stabilize this additional variability. Empirically, the Gaussian simulations in Section 4 favor relaxed SVEM fits, whereas the binomial simulations favor non-relaxed fits; we default to relaxed = TRUE for Gaussian and relaxed = FALSE for binomial responses. B.3 provides additional diagnostic summaries of selected $(\lambda,\gamma)$ distributions and bias–variance behavior in representative Gaussian and binomial settings.
@@ 2.3.1 Deterministic expansion utilities
A practical convenience in SVEMnet is that the user lists main-effect predictors only once and lets the expansion utilities (bigexp_terms() and bigexp_formula()) deterministically build interaction and polynomial columns. The same specification can be reused across responses, guaranteeing identical design matrices, stable column order, and reproducible comparisons. With factorial_order = 2 and polynomial_order = 2, for example, the generated right-hand side includes all main effects, two-way interactions, and quadratic terms (and, optionally, partial-cubic terms of the form $\mathrm{I}(X^{2})\!:\!Z$) without hand-writing products or powers; categorical factors are handled using the contrasts recorded for the training data. We also include any mixture factors in the full expansion rather than constructing Scheffé polynomials [38].
When blocking or covariate factors (e.g., Operator, day, plate, ambient temperature) are supplied via blocking in bigexp_terms(), these factors make an additive adjustment to the response surface and do not interact with the other study factors.
Discrete numeric predictors
Some process variables are recorded as numeric but are only feasible to run at a discrete set of levels (e.g., fixed pump settings or allowable N:P ratios). To avoid proposing unattainable intermediate settings in candidate generation and whole-model diagnostics, SVEMnet allows users to mark such predictors via the discrete_numeric argument in bigexp_terms() (which only affects the candidate recipe settings, not the SVEM model fit). The resulting specification records the allowed numeric support in the training data, and downstream tools (random candidate generation and WMT evaluation-set construction) restrict sampling to these observed levels.
Practical guidance on choosing expansion order
When the true response surface order is unknown, we recommend starting with a moderately rich expansion that is scientifically plausible (e.g., third-order polynomials with up to three-way interactions and partial cubic terms for mixture–process screening), and allowing the elastic-net SVEM fit to downselect. This “start wide, shrink down” strategy is consistent with the SVEM-based LNP workflow of Karl et al. [19] and is supported by the simulation results in Section 4, where mild over-specification is typically less harmful than gross under-specification. In practice we recommend: (i) fit a candidate expansion; (ii) examine actual-by-predicted diagnostics and predictive uncertainty; (iii) optionally apply WMT (Section 2.5) to assess global signal; and (iv) if results are unstable, compare two adjacent expansion orders (e.g., Orders 2 and 3) and check whether the top candidate sets and fitted predictions are materially changed. A simple candidate stability check is to re-run scoring under both expansions and report the overlap of top-ranked recipes and medoids (B.6).
@@ 2.4 Selection criterion
Figure 1: Within-replicate SVEMnet tuning flow for one bootstrap member $b$.
@@
Heuristic status and boundary behavior
Because the validation copy is itself random under FRW resampling, wAIC and wBIC are best interpreted as validation-weighted tuning scores rather than as classical information criteria with formal asymptotic guarantees. When $w_{i}^{\mathrm{valid}}\equiv 1$ these criteria reduce (up to constants independent of $\lambda$) to the standard Gaussian AIC/BIC forms, and the use of $n_{\mathrm{eff}}^{\mathrm{adm}}$ in the BIC-style penalty can be viewed as a simple effective-size adjustment for highly variable validation weights. Empirically, the primary motivation for wAIC and wBIC in SVEM is stability near the interpolation boundary $n\approx p_{\mathrm{full}}$. B.1 provides a focused simulation around $n=p_{\mathrm{full}}$ that compares wAIC/wBIC with the loss-only wSSE selector in terms of selected model size and out-of-sample error.
2.5 Permutation-based whole-model diagnostic (WMT)
SVEMnet includes a permutation-based whole-model test (WMT) heuristic for Gaussian SVEM fits [5]. The goal is to provide a practical diagnostic for whether a fitted SVEM model exhibits global factor signal beyond what would be expected under a constant-response null, complementing the usual actual-by-predicted and residual plots. WMT is a randomized, simulation-based procedure and the resulting “$p$-values” are approximate; we treat them as diagnostic summaries rather than as formal hypothesis tests.
At a high level, WMT compares standardized SVEM predictions on a space-filling set of evaluation points with a reference distribution obtained by refitting the same SVEM specification to permuted responses. Let $T=\{x_{i}\}_{i=1}^{n_{\mathrm{point}}}$ be a space-filling set over the modeled factor space (respecting mixture constraints, categorical levels, and any recorded discrete numeric supports). For each evaluation point $x\in T$, we compute a standardized prediction
$h(x)=\frac{\hat{f}(x)-\bar{y}}{\hat{s}(x)},$ (2)
where $\hat{f}(x)$ is the SVEM ensemble mean prediction, $\bar{y}$ is the training-set mean response, and $\hat{s}(x)$ is the ensemble standard deviation across SVEM bootstrap predictions at $x$ (a pointwise uncertainty proxy) [5]. Repeating this calculation for $n_{\mathrm{SVEM}}$ independent SVEM refits (new FRW weights each time) yields an $n_{\mathrm{SVEM}}\times n_{\mathrm{point}}$ matrix of standardized predictions. WMT then fits the same SVEM specification to $n_{\mathrm{perm}}$ random permutations of the response vector and records the corresponding standardized predictions on $T$ in an $n_{\mathrm{perm}}\times n_{\mathrm{point}}$ reference matrix.
To summarize global deviation from the null, WMT standardizes the permutation matrix columnwise, applies a reduced-rank singular value decomposition (retaining the leading components explaining a user-specified fraction of the permutation variance), and computes Mahalanobis-like distances for the permutation rows and for the original-data refits, following Karl [5]. The resulting distance distributions are displayed graphically; clear separation indicates strong global signal. For a scalar summary, SVEMnet fits a flexible parametric distribution to the permutation distances (SHASH [39]) and reports the fitted right-tail probability as a whole-model “$p$-value.” Because the procedure uses finite budgets ($n_{\mathrm{perm}}$, $n_{\mathrm{point}}$, $n_{\mathrm{SVEM}}$), results have Monte Carlo variability; for interpretability, users should set and report random seeds and treat marginal cases cautiously. The current SVEMnet defaults are $n_{\mathrm{point}}=2000$, $n_{\mathrm{SVEM}}=10$, $n_{\mathrm{perm}}=150$, and a 90% variance-retention threshold for the reduced-rank distance calculation.
In SVEMnet, WMT can be run for multiple responses via svem_wmt_multi(), which reuses a shared deterministic expansion and constraint specification. The resulting WMT “multipliers” (monotone transforms of the per-response approximate $p$-values) can optionally reweight user-specified multi-response desirability weights when producing the diagnostic wmt_score column (Section 2.6). In this paper we treat WMT reweighting as a secondary sensitivity check rather than as the primary decision rule, consistent with the diagnostic intent of the test heuristic.
@@ 2.6.1 Sampling and specification limits
Candidate settings are generated by svem_random_table_multi() and, in typical use, svem_score_random() calls this generator and then appends SVEM-based predictions, desirability scores, uncertainty summaries, and (optionally) mean-in-specification proxies. When specification limits are provided for a response, SVEMnet uses the SVEM bootstrap to estimate, at each candidate, the probability that the process mean lies inside those limits and adds per-response and joint mean-level specification probabilities (e.g., prob_in_spec and p_joint_mean) as additional columns in the scored table. These mean-level probabilities are used only as ranking proxies: formal specification limits apply to individual observations, not to the process mean.
@@ 2.6.2 Scoring and response weighting
$S(x)=\exp\!\left(\sum_{r}w_{r}\log\big(d_{r,\varepsilon}(x)\big)\right),\qquad d_{r,\varepsilon}(x)=(1-\varepsilon)\,d_{r}(x)+\varepsilon.$ (3)
@@
Practical desirability calibration and sensitivity checks
In practice we recommend choosing Derringer–Suich parameters from scientific acceptability ranges rather than from the fitted model itself. For each response, the user specifies a goal (maximize, minimize, or hit a target) together with bounds and optionally a target. The bounds define where desirability is essentially zero (clearly unacceptable) and where it saturates at one (fully acceptable or “good enough”); the shape exponent controls how rapidly desirability approaches one between these values. SVEMnet uses a simple default shape (exponent =1=1) unless the user supplies custom exponents, which yields a linear ramp on the response scale. Exponents greater than one make the desirability curve more stringent near the target, while exponents below one make it more tolerant.
The weighted geometric mean in (3) is used because it is the standard aggregation in the desirability literature and because it enforces “weakest-link” behavior: if any key response has near-zero desirability, the overall score is driven toward zero, reflecting the fact that a single out-of-spec response can invalidate a recipe [41]. Alternatives such as arithmetic means can mask poor performance in a critical response.
When desirability parameters encode subjective trade-offs, simple sensitivity checks are valuable. A practical approach is to rerun scoring under plausible alternative bounds, targets, exponents, or response weights and then compare the overlap and rank stability of the top candidates and medoids. B.5 provides a compact desirability-calibration sensitivity check.
User-specified goal weights $\tilde{w}_{r}$ are first normalized to $\bar{w}_{r}$ with $\sum_{r}\bar{w}_{r}=1$. When reweight_by_wmt = FALSE, we set $w_{r}=\bar{w}_{r}$ and report a single column score based on $S(x)$. When reweight_by_wmt = TRUE for Gaussian responses, we further reweight the normalized user weights by a monotone function of the permutation whole-model $p$-values from [5] (for example $\log_{10}(p_{r})$), renormalize to obtain WMT-adjusted weights $w_{r}^{\mathrm{WMT}}$, and compute a second geometric-mean score wmt_score by using $w_{r}=w_{r}^{\mathrm{WMT}}$ in $S(x)$. These WMT-adjusted weights are diagnostic and do not alter the underlying SVEM fits. In this paper we treat the resulting WMT-based multipliers as a secondary priority: the primary optimization uses the user-specified weights $\tilde{w}_{r}$, and wmt_score is intended to provide additional, lower-priority optimal candidates when budget allows.
@@
For candidate scoring, blocking variables are held fixed at a single reference setting while the controllable factors are varied over the feasible region. Categorical blocking factors are evaluated at the most common observed level, and continuous blocking covariates are evaluated at the midpoint of their modeled range. Under the additive-only blocking assumption this defines a consistent reference for comparing controllable settings; if desirability bounds/targets are specified on an absolute response scale, users may optionally re-score at alternative blocking levels to confirm stability of the top candidates.